OpenStudy (anonymous):

Solve for x 5x+3=x-7 Will medal and fan

OpenStudy (anonymous):

@Mashy

OpenStudy (anonymous):

@matt101

OpenStudy (matt101):

You need to isolate x here. What would be your first step?

OpenStudy (anonymous):

adding x to both sides

OpenStudy (anonymous):

is that correct ? @matt101

OpenStudy (matt101):

Well let's see. If you add x to both sides, you end up with 6x+3=2x-7. We're not really any closer to isolating x. What can we do to your original equation to get all the x's on the same side?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

does 6x=-10

OpenStudy (anonymous):

and x =-10/6 in fraction form

OpenStudy (matt101):

Close! How did you get rid of the x on the right side?

OpenStudy (anonymous):

well I did mathematical operations on both sides

OpenStudy (anonymous):

so does x= -10/6 in fraction form

OpenStudy (anonymous):

@matt101

OpenStudy (matt101):

You added x to the left, but subtracted it from the right. You need to do the same operation to both sides in order to keep the sides of the equation equal. In this case, you need to subtract x from both sides. When you do that, you eliminate the x on the right: \[5x+3=x-7\]\[5x+3-x=x-7-x\]\[4x+3=-7\] Does that make sense? Can you take it from there?

OpenStudy (anonymous):

why is it 4x shouldnt it be 6x

OpenStudy (matt101):

What's 5x-x?

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@dan815